# Artin-Schreier towers of finite fields

### v1

@@ Abstract.
Given a prime number $p$, we consider the tower of finite fields $F_{p}=L_{-1}\subset L_{0}\subset L_{1}\subset\cdots$, where each step corresponds to an Artin-Schreier extension of degree $p$, so that for $i\geq 0$, $L_{i}=L_{i-1}[c_{i}]$, where $c_{i}$ is a root of $X^{p}-X-a_{i-1}$ and $a_{i-1}=(c_{-1}\cdots c_{i-1})^{p-1}$, with $c_{-1}=1$. We extend and strengthen to arbitrary primes prior work of Popovych for $p=2$ on the multiplicative order of the given generator $c_{i}$ for $L_{i}$ over $L_{i-1}$. In particular, for $i\geq 0$, we show that $O(c_{i})=O(a_{i})$, except only when $p=2$ and $i=1$, and that $O(c_{i})$ is equal to the product of the orders of $c_{j}$ modulo $L_{j-1}^{\times}$, where $0\leq j\leq i$ if $p$ is odd, and $i\geq 2$ and $1\leq j\leq i$ if $p=2$. We also show that for $i\geq 0$, the $\mathrm{Gal}(L_{i}/L_{i-1})$-conjugates of $a_{i}$ form a normal basis of $L_{i}$ over $L_{i-1}$. In addition, we obtain the minimal polynomial of $c_{1}$ over $F_{p}$ in explicit form.
@@ 1. Introduction
We fix throughout the paper a prime number $p$ and an algebraic closure $E$ of the field $F_{p}$ with $p$ elements. Recursively define $c_{-1},c_{0},c_{1},\dots\in E$ by taking $c_{-1}=1$ and selecting $c_{i+1}$ so that
@@
(1.1) $L_{-1}\subset L_{1}\subset L_{2}\subset\cdots,$
@@
The first extension $L_{0}/F_{p}$ is connected to the period conjecture for the Bell numbers modulo $p$, which reduces to the question of whether or not the root $c_{0}$ of $X^{p}-X-1$ has multiplicative order $N_{0}$. This problem has been addressed, among others, by Wagstaff [Wa]; Car, Gallardo, Rahavandrainy, and Vaserstein [CGRV]; Montgomery, Nahm, and Wagstaff [MNW]; and Gallardo in [G1] and [G2]. While significant progress has been made, it remains open, as well as the question of whether $N_{0}$ is always square-free.
@@
While finding the exact orders of given elements in a finite field may be too difficult, it is sometimes possible to of find lower bounds for the orders of these elements. There is a vast literature on the subject and we refer the reader to the paper [BR] by Brochero Martínez and Reis, and the article [DMP] by Dose, Mercuri, and Pal, as well as references therein.
Albert’s construction was rediscovered by Ito, Kashiwara, and Huling [IKH]. They give an algorithm to produce irreducible polynomials of high degree related to the tower (1.1), and for $p=2$ they discuss the problem of finding primitive elements for each member of this tower.
@@
One of the most recent papers studying the tower (1.1) in the case $p=2$ is due to Popovych [P]. He gives a lower bound for the order of $c_{i}$ as well as sufficient conditions for $O(c_{i})$, $i\geq 2$, to attain its largest possible value, which he proves to be $N_{1}\cdots N_{i}$. Under these conditions, he shows that one also has $O(a_{i})=O(c_{i})$, $i\geq 2$. As $O(c_{0})=3$ and the Fermat numbers $N_{i}$, $i\geq 1$, are relatively prime to $N_{0}=3$, it follows from (1.2) that for $i\geq 2$, $O(c_{i})=N_{1}\cdots N_{i}$ if and only if $c_{0}c_{i}$ is a primitive element of $L_{i}$. Questions of a similar nature had been addressed earlier by Wiedemann [W], who found other potential primitive elements for $L_{i}$. However, his construction of the fields in the tower (1.1) was different, and an explicit relationship between the two constructions was left open. Seyfarth and Ranade [SR] show how to translate the problem of the primitivity of Wiedemann’s elements into a specific feature of the characteristic polynomial of certain matrices over $F_{2}$.
@@
In this paper we extend Popovych’s results from [P] to all primes. More explicitly, Theorem 4.5 extends and strengthens [P, Theorem 4] by showing, unconditionally, that $O(c_{i})=O(a_{i})$ for all $i\geq 0$, except only when $p=2$ and $i=1$. In Theorem 4.6, which extends [P, Theorem 3], we give a formula for the common order of $c_{i}$ and $a_{i}$ as a product of the orders of $c_{j}$ modulo $L_{j-1}^{\times}$, where $0\leq j\leq i$ if $p$ is odd, and $i\geq 2$ and $1\leq j\leq i$ if $p=2$. Corollary 4.7 provides necessary and sufficient conditions for $O(c_{i})$ to attain its largest possible value, namely $N_{0}\cdots N_{i}$ if $p$ is odd, and $N_{1}\cdots N_{i}$ if $p=2$ and $i\geq 2$. As $p-1$ is relatively prime to all $N_{i}$ (see Lemma 3.1), it follows from (1.2) that if $p$ is odd, $i\geq 0$, and $z$ is a primitive element of $F_{p}$, then $o(c_{i})=N_{0}\cdots N_{i}$ if and only if $zc_{i}$ is a primitive element of $L_{i}$. Corollary 4.8 extends [P, Corollary 2] and gives a lower bound for the order of $c_{i}$.
Moreover, in Section 5 we show that for all $i\geq 0$, the $\mathrm{Gal}(L_{i}/L_{i-1})$-conjugates of $a_{i}$ form a normal basis of $L_{i}$ over $L_{i-1}$.
Furthermore, in Section 6 we establish a connection between the minimal polynomials of $c_{i}$ and $a_{i-1}$ over $F_{p}$, $i\geq 1$, which is used in Theorem 6.5 to explicitly exhibit the minimal polynomial of $c_{1}$ over $F_{p}$.
We begin the paper by reviewing Albert’s construction of Artin-Schreier field towers in Section 2 and their basic properties in Section 3.
@@ 2. Construction of the towers
$\left(\begin{array}[]{cccccc}0&a&0&\cdots&0&0\\ 0&1&a&\cdots&0&0\\ 0&0&1&\cdots&0&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&0&\cdots&1&a\\ 1&0&0&\cdots&0&1\\ \end{array}\right).$
@@
Adding them up and appealing to (2.1), we obtain
@@
It should be noted that that union, say $L$, of the members of the tower
@@
is the smallest subfield of $E$ that has no extensions of degree $p$. By construction, a basis for $L$ over $F_{p}$ is formed by 1 and all elements of the form $c_{n_{1}}^{e_{1}}\cdots c_{n_{k}}^{e_{k}}$, where $k\geq 1$, $0\leq n_{1}<\dots<n_{k}$ and $1\leq e_{1},\dots,e_{k}<p$. We can also construct $L$ as the quotient of the polynomial ring $F_{p}[X_{-1},X_{0},X_{1},\dots]$ by the ideal $I$ generated by
@@
Indeed, there is a ring epimorphism $F_{p}[X_{-1},X_{0},X_{1},\dots]\to L$ mapping $X_{j}\to c_{j}$ and containing (2.7) in its kernel. The resulting epimporphism from the quotient ring $F_{p}[X_{-1},X_{0},X_{1},X_{2},\dots]/I$ to $L$ is seen to be injective by means of the preceding basis of $L$.
@@ 3. Basic properties
If $i\neq j$ then $N_{i}$ and $N_{j}$ are relatively prime.
@@
Since $c_{-1}\dots c_{i}\in L_{i}$, (2.5) gives $(c_{-1}\dots c_{i})^{p^{p^{i+1}}-1}=1.$ But $N_{0}\cdots N_{i}=\frac{p^{p^{i+1}}-1}{p-1}$ and $a_{i}=(c_{-1}\dots c_{i})^{p-1}$, so $a_{i}^{N_{0}\cdots N_{i}}=1$. If $i=0$ then Lemma 3.4 yields $c_{0}^{N_{0}}=a_{-1}=1$ as well as $c_{i+1}^{N_{0}\cdots N_{i}N_{i+1}}=a_{i}^{N_{0}\cdots N_{i}}=1.$
When $p=2$ a sharpening of Lemma 3.4 is possible.
@@
Let $F\subset K$ be a tower of fields. Suppose $a\in K^{\times}$ has finite order and set $m=\min\{n\geq 1\,|\,a^{n}\in L^{\times}\}$. Then $O(a)=m\times O(a^{m})$.
Proof.
Set $G=K^{\times}$ and consider its subgroups $N=L^{\times}$ and $H=\langle a\rangle$. Then
$|H|=|H/H\cap N|\times|H\cap N|=m\times O(a^{m}).\qed$
@@ 4. The orders of $c_{i}$ and $a_{i}$
Suppose, if possible, that $q$ is a prime factor of both numbers. Then $q|N_{j}$ for some $0\leq j\leq i$. If $j<i$ then $q|(p^{p^{i}}-1)$. But $q|(N_{i}-(p^{p^{i}}-1))$, so $q|N_{i}$, against Corollary 3.3. If $j=i$ then from $q|(N_{i}-(p^{p^{i}}-1))$ we infer $q|(p^{p^{i}}-1)$, against Lemma 3.1. ∎
@@
It follows from Lemma 4.3, (4.2) and (4.3) that
@@
If $p$ is odd, the result follows immediately by repeatedly applying Proposition 4.4 and Theorem 4.5. Suppose $p=2$. That $O(c_{0})=N_{0}$ and $O(c_{1})=N_{0}N_{1}$ was shown in Lemma 3.7. As $O(a_{1})=N_{1}$ by Lemma 3.7, we infer $O(c_{2})=N_{2}N_{1}$ by Proposition 4.4 and the fact that $M_{2}=N_{2}$. Repeated application of Proposition 4.4 and Theorem 4.5 now give $O(c_{i})=N_{1}N_{2}M_{3}\cdots M_{i}$ for all $i\geq 3$. As $M_{3}=N_{3}$ and $M_{4}=N_{4}$, the proof is complete. ∎
@@ 5. A normal basis for $L_{i}$ over $L_{i-1}$
Here we show that the Galois conjugates of $a_{i}$ form a normal basis for $L_{i}$ over $L_{i-1}$ for all $i\geq 1$.
@@
Thus $M=DVS$, where $D=\mathrm{diag}(\binom{p-1}{0},\binom{p-1}{1},\dots,\binom{p-1}{p-1})$, $V$ is the Vandermonde matrix $V(0\times(-1)^{i},1\times(-1)^{i},\dots,(p-1)\times(-1)^{i})$, and $S=\mathrm{diag}(a_{i-1},\dots,a_{i-1})$. ∎
@@ 6. Minimal polynomials
$f=\underset{h\in Gg}{\Pi}h.$
In particular, $\deg f=\deg h\times|G|/|G_{g}|$.
@@
If $i\geq 0$, then $c_{i}$ and $a_{i}$ have degree $p^{i+1}$ over $F_{p}$.
@@
By Theorem 6.2 the degree of $c_{i}$ over $F_{p}$ is a factor of $p^{i+1}$. If the degree is not $p^{i+1}$, then necessarily $c_{i}^{p^{p^{i}}}=c_{i}$, which contradicts Lemma 2.5. Likewise, if the degree of $a_{i}$ is not $p^{i+1}$, then Lemma 2.5 yields
@@
which implies $c_{i}^{p-1}=(c_{i}+(-1)^{i})^{p-1}$. It follows that $(c_{i}+(-1)^{i})c_{i}^{-1}\in F_{p}$. But $(c_{i}+(-1)^{i})c_{i}^{-1}=1+(-1)^{i}c_{i}^{-1}$, which is not in $F_{p}$ by (2.4). ∎
@@
$f_{c_{i},F_{p}}(X)=f_{a_{i-1},F_{p}}(X^{p}-X).$
In particular, this polynomial is irreducible in $F_{p}[X]$.
@@
By definition, $f_{c_{i},L_{i-1}}(X)=X^{p}-X-a_{i-1}$. By Theorem 6.3, $a_{i-1}$ has exactly $p^{p^{i}}$ conjugates over $F_{p}$, so Theorem 6.1 yields
$f_{c_{i},F_{p}}(X)=(X^{p}-X-a_{i-1})(X^{p}-X-a_{i-1}^{p})\cdots(X^{p}-X-a_{i-1}^{p^{p^{i}-1}})=f_{{a_{i-1},F_{p}}}(X^{p}-X).\qed$
@@
The minimal polynomial of $c_{1}$ over $F_{p}$ is equal to
@@
In particular, this polynomial is irreducible in $F_{p}[X]$.
@@
We have $f_{c_{0},F_{p}}(X)=X^{p}-X-1$, so $f_{c_{0}^{p-1},F_{p}}(X)=X^{p}-(X-1)^{p-1}-2$ by Lemma 2.2. Thus, by Theorem 6.4,
$f_{c_{1},F_{p}}(X)$ $=f_{c_{0}^{p-1},F_{p}}(X^{p}-X)=(X^{p}-X)^{p}-(X^{p}-X-1)^{p-1}-2$

### v2

@@ Abstract.
Given a prime number $p$, we consider the tower of finite fields $\mathbb{F}_{p}=L_{-1}\subset L_{0}\subset L_{1}\subset\cdots$, where each step corresponds to an Artin-Schreier extension of degree $p$, so that for $i\geq 0$, $L_{i}=L_{i-1}[c_{i}]$, where $c_{i}$ is a root of $X^{p}-X-a_{i-1}$ and $a_{i-1}=(c_{-1}\cdots c_{i-1})^{p-1}$, with $c_{-1}=1$. We extend and strengthen to arbitrary primes prior work of Popovych for $p=2$ on the multiplicative order $O(c_{i})$ of the given generator $c_{i}$ for $L_{i}$ over $L_{i-1}$. In particular, for $i\geq 0$, we show that $O(c_{i})=O(a_{i})$, except only when $p=2$ and $i=1$, and that $O(c_{i})$ is equal to the product of the orders of $c_{j}$ modulo $L_{j-1}^{\times}$, where $0\leq j\leq i$ if $p$ is odd, and $i\geq 2$ and $1\leq j\leq i$ if $p=2$. We also show that for $i\geq 0$, the $\mathrm{Gal}(L_{i}/L_{i-1})$-conjugates of $a_{i}$ form a normal basis of $L_{i}$ over $L_{i-1}$. In addition, we obtain the minimal polynomial of $c_{1}$ over $\mathbb{F}_{p}$ in explicit form.
@@ 1. Introduction
We fix throughout the paper a prime number $p$ and an algebraic closure $E$ of the field $\mathbb{F}_{p}$ with $p$ elements. Recursively define $c_{-1},c_{0},c_{1},\dots\in E$ by taking $c_{-1}=1$ and selecting $c_{i+1}$ so that
@@
(1.1) $L_{-1}\subset L_{0}\subset L_{1}\subset L_{2}\subset\cdots,$
@@
The first extension $L_{0}/\mathbb{F}_{p}$ is connected to the period conjecture for the Bell numbers modulo $p$, which reduces to the question of whether or not the root $c_{0}$ of $X^{p}-X-1$ has multiplicative order $N_{0}$. This problem has been addressed, among others, by Wagstaff [Wa]; Car, Gallardo, Rahavandrainy, and Vaserstein [CGRV]; Montgomery, Nahm, and Wagstaff [MNW]; and Gallardo in [G1] and [G2]. While significant progress has been made, it remains open, as well as the question of whether $N_{0}$ is always square-free.
@@
Albert’s construction was rediscovered by Ito, Kashiwara, and Huling [IKH]. They give an algorithm to produce irreducible polynomials of high degree related to the tower (1.1), and for $p=2$ they discuss the problem of finding primitive elements for each member of this tower. While finding the exact orders of given elements in a finite field may be too difficult, it is sometimes possible to find lower bounds for the orders of these elements. There is a vast literature on the subject and we refer the reader to the paper [BR] by Brochero Martínez and Reis, and the article [DMPS] by Dose, Mercuri, Pal, and Stirpe, as well as references therein.
@@
One of the most recent papers studying the tower (1.1) in the case $p=2$ is due to Popovych [P]. He gives a lower bound for the order $O(c_{i})$ of $c_{i}$ as well as sufficient conditions for $O(c_{i})$, $i\geq 2$, to attain its largest possible value, which he proves to be $N_{1}\cdots N_{i}$. Under these conditions, he shows that one also has $O(a_{i})=O(c_{i})$, $i\geq 2$. As $O(c_{0})=3$ and the Fermat numbers $N_{i}$, $i\geq 1$, are relatively prime to $N_{0}=3$, it follows from (1.2) that for $i\geq 2$, $O(c_{i})=N_{1}\cdots N_{i}$ if and only if $c_{0}c_{i}$ is a primitive element of $L_{i}$. Questions of a similar nature had been addressed earlier by Wiedemann [W], who found other potentially primitive elements for $L_{i}$. However, his construction of the fields in the tower (1.1) was different, and an explicit relationship between the two constructions was left open. Seyfarth and Ranade [SR] show how to translate the problem of the primitivity of Wiedemann’s elements into a specific feature of the characteristic polynomial of certain matrices over $F_{2}$. Recursive constructions of elements of predictably high order in finite fields are useful for the development of discrete logarithm ciphers and pseudo-random number generators, and there are added benefits if the elements of high order determine normal bases over a ground field (see [PS]).
@@
In this paper we extend Popovych’s results from [P] to all primes. In particular, Theorem 4.5 extends and strengthens [P, Theorem 4] by showing, unconditionally, that $O(c_{i})=O(a_{i})$ for all $i\geq 0$, except only when $p=2$ and $i=1$. In Corollary 4.7, which extends [P, Theorem 3], we give a formula for the common order of $c_{i}$ and $a_{i}$ as a product of the orders of $c_{j}$ modulo $L_{j-1}^{\times}$, where $0\leq j\leq i$ if $p$ is odd, and $i\geq 2$ and $1\leq j\leq i$ if $p=2$. Corollary 4.8 provides necessary and sufficient conditions for $O(c_{i})$ to attain its largest possible value, namely $N_{0}\cdots N_{i}$ if $p$ is odd, and $N_{1}\cdots N_{i}$ if $p=2$ and $i\geq 2$. As $p-1$ is relatively prime to all $N_{i}$ (see Lemma 3.1), it follows from (1.2) that if $p$ is odd, $i\geq 0$, and $z$ is a primitive element of $\mathbb{F}_{p}$, then $O(c_{i})=N_{0}\cdots N_{i}$ if and only if $zc_{i}$ is a primitive element of $L_{i}$. Corollary 4.9 extends [P, Corollary 2] and gives a lower bound for the order of $c_{i}$.
We begin the paper by reviewing Albert’s construction of Artin-Schreier field towers in Section 2 and establishing their basic properties in Section 3. Section 4 is devoted to the proof of Theorem 4.5 and its consequences. In Section 5 and Section 6 we establish two further properties of interest. Theorem 5.1 shows that for all $i\geq 0$, the $\mathrm{Gal}(L_{i}/L_{i-1})$-conjugates of $a_{i}$ form a normal basis of $L_{i}$ over $L_{i-1}$. Theorem 6.4 establishes a connection between the minimal polynomials of $c_{i}$ and $a_{i-1}$ over $\mathbb{F}_{p}$, $i\geq 1$, which is applied in Corollary 6.5 to explicitly exhibit the minimal polynomial of $c_{1}$ over $\mathbb{F}_{p}$.
@@ 2. Construction of the towers
$\begin{bmatrix}0&a&0&\cdots&0&0\\ 0&1&a&\cdots&0&0\\ 0&0&1&\cdots&0&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&0&\cdots&1&a\\ 1&0&0&\cdots&0&1\\ \end{bmatrix}.$
@@
Adding them up and appealing to Lemma 2.4, we obtain
@@
It should be noted that the union, say $L$, of the members of the tower
@@
is the smallest subfield of $E$ that has no extensions of degree $p$. By construction, a basis for $L$ over $\mathbb{F}_{p}$ is formed by 1 and all elements of the form $c_{n_{1}}^{e_{1}}\cdots c_{n_{k}}^{e_{k}}$, where $k\geq 1$, $0\leq n_{1}<\dots<n_{k}$ and $1\leq e_{1},\dots,e_{k}<p$. We can also construct $L$ as the quotient of the polynomial ring $\mathbb{F}_{p}[X_{-1},X_{0},X_{1},\dots]$ by the ideal $I$ generated by
@@
Indeed, there is a ring epimorphism $\mathbb{F}_{p}[X_{-1},X_{0},X_{1},\dots]\to L$ mapping $X_{j}\to c_{j}$ and containing (2.5) in its kernel. The resulting epimorphism from the quotient ring $\mathbb{F}_{p}[X_{-1},X_{0},X_{1},X_{2},\dots]/I$ to $L$ is seen to be injective by means of the preceding basis of $L$.
@@ 3. Basic properties
In this section we will establish some basic properties of the sequence of numbers $N_{i}$, $i\geq 0$, for a fixed prime $p$. These properties were noticed earlier by Popovych [P] in the case $p=2$; for odd primes the proofs are similar.
@@
If $i\neq j$, then $N_{i}$ and $N_{j}$ are relatively prime.
@@
Since $c_{-1}\dots c_{i}\in L_{i}$, (2.4) gives $(c_{-1}\dots c_{i})^{p^{p^{i+1}}-1}=1.$ But $N_{0}\cdots N_{i}=\frac{p^{p^{i+1}}-1}{p-1}$ and $a_{i}=(c_{-1}\dots c_{i})^{p-1}$, so $a_{i}^{N_{0}\cdots N_{i}}=1$. If $i=0$, then Lemma 3.4 yields $c_{0}^{N_{0}}=a_{-1}=1$ as well as $c_{i+1}^{N_{0}\cdots N_{i}N_{i+1}}=a_{i}^{N_{0}\cdots N_{i}}=1.$
When $p=2$ a sharpening of Lemma 3.4 is possible. The next result follows from the fact that finite subgroups of the multiplicative group of a field are cyclic.
@@
Let $K_{1}\subset K_{2}$ be an extension of fields. Suppose $a\in K_{2}^{\times}$ has finite order and set $m=\min\{n\geq 1\,|\,a^{n}\in K_{1}^{\times}\}$. Then $O(a)=m\times O(a^{m})$.
@@ 4. The orders of $c_{i}$ and $a_{i}$
Suppose, if possible, that $q$ is a prime factor of both numbers. Then $q|N_{j}$ for some $0\leq j\leq i$. If $j<i$, then $q|(p^{p^{i}}-1)$. But $q|(N_{i}-(p^{p^{i}}-1))$, so $q|N_{i}$, against Corollary 3.3. If $j=i$, then from $q|(N_{i}-(p^{p^{i}}-1))$ we infer $q|(p^{p^{i}}-1)$, against Lemma 3.1. ∎
@@
It follows from (4.2), (4.3), and Lemma 4.3 that
@@
If $p$ is odd, the result follows immediately by repeatedly applying Proposition 4.4 and Theorem 4.5. If $p=2$, the result follows from Remark 4.6 and a repeated application of Proposition 4.4 and Theorem 4.5. ∎
@@ 5. A normal basis for $L_{i}$ over $L_{i-1}$
Here we show that the Galois conjugates of $a_{i}$ form a normal basis for $L_{i}$ over $L_{i-1}$ for all $i\geq 1$. Given a field $K$ and $\alpha_{1},\dots,\alpha_{n}\in K$, we write $V(\alpha_{1},\dots,\alpha_{n})$ for the Vandormonde matrix in $M_{n}(K)$ associated to $(\alpha_{1},\dots,\alpha_{n})$, that is, with rows equal to $(1,\dots,1),(\alpha_{1},\dots,\alpha_{n}),\dots,(\alpha_{1}^{n-1},\dots,\alpha_{n}^{n-1})$.
@@
Thus $M=DVS$, where $D=\mathrm{diag}(\binom{p-1}{0},\binom{p-1}{1},\dots,\binom{p-1}{p-1})$, $V$ is the Vandermonde matrix $V(0\times(-1)^{i},1\times(-1)^{i},\dots,(p-1)\times(-1)^{i})$, and $S=\mathrm{diag}(a_{i-1},\dots,a_{i-1})$.
@@ 6. Minimal polynomials
$f=\underset{h\in Gg}{\Pi}h,$
where $Gg$ is the orbit of $g$ under the action of $G$. In particular, we have that $\deg f=\deg h\times|G|/|G_{g}|$, where $G_{g}$ is the stabilizer of $g$ in $G$.
@@
If $i\geq 0$, then $c_{i}$ and $a_{i}$ have degree $p^{i+1}$ over $\mathbb{F}_{p}$.
@@
By Theorem 6.2 the degree of $c_{i}$ over $\mathbb{F}_{p}$ is a factor of $p^{i+1}$. If the degree is not $p^{i+1}$, then necessarily $c_{i}^{p^{p^{i}}}=c_{i}$, which contradicts Lemma 2.5. Likewise, if the degree of $a_{i}$ is not $p^{i+1}$, then Lemma 2.5 yields
@@
which implies $c_{i}^{p-1}=(c_{i}+(-1)^{i})^{p-1}$. It follows that $(c_{i}+(-1)^{i})c_{i}^{-1}\in\mathbb{F}_{p}$. But $(c_{i}+(-1)^{i})c_{i}^{-1}=1+(-1)^{i}c_{i}^{-1}$, which is not in $\mathbb{F}_{p}$ by (2.3). ∎
@@
$f_{c_{i},\mathbb{F}_{p}}(X)=f_{a_{i-1},\mathbb{F}_{p}}(X^{p}-X).$
In particular, this polynomial is irreducible in $\mathbb{F}_{p}[X]$.
@@
By definition, $f_{c_{i},L_{i-1}}(X)=X^{p}-X-a_{i-1}$. By Theorem 6.3, $a_{i-1}$ has exactly $p^{p^{i}}$ conjugates over $\mathbb{F}_{p}$, so Theorem 6.1 yields
$f_{c_{i},\mathbb{F}_{p}}(X)=(X^{p}-X-a_{i-1})(X^{p}-X-a_{i-1}^{p})\cdots(X^{p}-X-a_{i-1}^{p^{p^{i}-1}})=f_{{a_{i-1},\mathbb{F}_{p}}}(X^{p}-X).\qed$
@@
The minimal polynomial of $c_{1}$ over $\mathbb{F}_{p}$ is equal to
@@
In particular, this polynomial is irreducible in $\mathbb{F}_{p}[X]$.
@@
We have $f_{c_{0},\mathbb{F}_{p}}(X)=X^{p}-X-1$, so $f_{c_{0}^{p-1},\mathbb{F}_{p}}(X)=X^{p}+(X-1)^{p-1}-2$ by Lemma 2.2. Thus, by Theorem 6.4,
$f_{c_{1},\mathbb{F}_{p}}(X)$ $=f_{c_{0}^{p-1},\mathbb{F}_{p}}(X^{p}-X)=(X^{p}-X)^{p}+(X^{p}-X-1)^{p-1}-2$